# Enumerating Complexity Revisited

### 07 Dec 2023 v1

@@ Abstract
We reduce the best-known upper bound on the length of a program that enumerates a set in terms of the probability of it being enumerated by a random program. We prove a general result that any linear upper bound for finite sets implies the same linear bound for infinite sets.
So far, the best-known upper bound was given by Solovay. He showed that the minimum length of a program enumerating a subset $S$ of natural numbers is bounded by minus three binary logarithms of the probability that a random program will enumerate $S$. Later, Vereshchagin showed that the constant can be improved from three to two for finite sets. In this work, using an improvement of the method proposed by Solovay, we demonstrate that any bound for finite sets implies the same for infinite sets, modulo logarithmic factors. Using Vereshchagin’s result, we improve the current best-known upper bound from three to two.
@@
In the first section, we introduce definitions of deterministic and randomized complexity, present the results known prior to our work, and state Theorem 1, which represents the main result of this work. In the second section, we associate cats with the deterministic machine and ants with the randomized machine, and we introduce shadow positions for the ants, which will follow the ants with some delay. We describe how the shadow positions will mimic the real ones. In the third and fourth sections, we demonstrate that if all accumulations in the shadow positions are serviced by cats, then all the necessary infinite sets will be enumerated. The fifth section contains a discussion of the results.
@@ 1 Introduction
To formulate the main results of this work, we will introduce the concepts of deterministic and randomized computing machines.
@@ 1.1 Deterministic Enumerating Machine
Let’s consider a deterministic machine, denoted as $D$, which takes a binary string as input and enumerates a certain subset of natural numbers. $D$ is not required to halt and may enumerate an infinite set. We define the complexity of a set, denoted as $I_{D}(S)$, as the minimum input length for $D$ to enumerate $S$. If there is no such input at all, we consider the complexity of the set to be infinite. As known, there exists a machine $D_{0}$ (such a machine is referred to as universal) in such a way that for any other machine $D,$ there is a constant $c$ such that:
@@
We can set $D$ equal to any such universal machine $D_{0}$ and define $I(S)$ as $I_{D}(S)$. Thus, $I(S)$ is defined up to a constant factor.
@@ 1.2 Probabilistic Enumerating Machine
Similar to the deterministic case, there exists a machine $M_{0}$ (referred to as optimal) in such a way that for any other $M,$ there is a constant $c$ such that:
@@
We can set $M$ equal to any such optimal machine $M_{0}$ and define $H(S)$ as $H_{M}(S).$ Therefore, $H(S)$ is defined up to a constant factor.
@@ 1.3 Connection Between Deterministic and Randomized Complexity
However, obtaining the converse inequality is much more challenging. Solovay proved a linear upper bound.
Theorem (Solovay, [4]). There exists a constant $c$ such that for any $S\subset\mathbb{N}$, it holds that $I(S)\leq 3\cdot H(S)+2\log H(S)+c.$
In his work, Vereshchagin improved the constant from three to two for finite sets:
Theorem (Vereshchagin, [5]). There exists a constant $c$ such that for any finite $S\subset\mathbb{N}$, it holds that $I(S)\leq 2\cdot H(S)+2\log H(S)+c.$
@@
Theorem 1. Let $\alpha\geq 1$ be such that for any finite $S\subset\mathbb{N}$,
@@
Applying this theorem to Vereshchagin’s theorem for finite sets, we improve the constant in Solovay’s bound from three to two (ignoring logarithmic factors).
To prove this theorem, we will introduce convenient terms. We will represent the behavior of deterministic and randomized machines as the movement of cats and ants on a directed graph.
@@ 2.6 Main Lemma
Lemma. Let $S$ be a set with a probability of enumeration greater than $\varepsilon$ (the threshold from the shift conditions). Let $S^{\prime}\subset S$ be a finite subset of it. Then at some point, a shift will occur for the set $X$ located between $S^{\prime}$ and $S$ (i.e., $S^{\prime}\subset X\subset S$).
Before proving this lemma, let’s show that it implies a desired bound for the complexity of enumerating infinite sets. Consider an increasing sequence of finite sets $S^{\prime}$ converging to $S$ and apply this lemma to each of them, obtaining, for each, its set $X$ (and one of the first $l(\varepsilon)$ cats servicing this set). Since there is a finite number of cats, one of them must have participated in servicing an infinite number of sets $S^{\prime}$, never leaving $S$ and being in the supersets of infinitely many sets $S^{\prime}$. Therefore, it will eventually enumerate $S$. Thus, we have:
@@
Now we will show how to use this to prove the main theorem.
Proof of Theorem 1. Let $k$ be an integer. We will show that if we set $\varepsilon=2^{-k}$ and $l(\varepsilon)=2^{\alpha k+O(\log k)}$, then during a shift, there will always be a cat among the first $l(\varepsilon)$ that comes to $X.$ Using $I(S)\leq\log_{2}l(\varepsilon),$ we will prove the theorem.
@@
Now, we need to prove the main lemma.
3 Proof of the Main Lemma
@@ 3.1 Number of Shifts
At every moment $t$ of the process, the Cantor space is divided into cones corresponding to ants. For each ant, we can count the total number of shifts in which it (itself or its ancestors) participated. We can define the set $W_{k}^{t}$ as the set in the Cantor space corresponding to ants with $k$ or more shifts by the time $t$. This is a basic (open and closed) set; the larger $k$ is, the smaller $W_{k}^{t}$ (for a given $t$). On the other hand, $W_{k}^{t}$ grows with increasing $t$. We can take the union over all $t$ and obtain an open set $W_{k}$ (it is even effectively open since the process is algorithmic - we assume that $\varepsilon$ is rational). We obtain a decreasing sequence of open sets:
@@ 3.2 The Case of Non-empty Intersection
In addition to $W_{\infty}$, consider the set $U$ in the Cantor space, consisting of those sequences for which the probabilistic machine $M$ enumerates the set $S$. Do the sets $U$ and $W_{\infty}$ intersect? We will show that if they do intersect, then the statement of the lemma is true, and then we will contradict the assumption that the intersection is empty.
@@ 4 Discussion
One of the most important questions in theoretical computer science is how much more efficient randomized computations are compared to deterministic ones. As Theorem 1 demonstrates, randomized computations are not superior to deterministic ones when considering infinite sets instead of finite ones. Moreover, the gap between complexities can be reduced to a factor of 2 (modulo logarithmic factors). To prove the theorem, we utilized a trick first used by Solovay: instead of cats (the deterministic machine) trying to catch the actual positions of ants (the randomized machine), the cats catch the shadow positions of the ants, which are slightly modified real positions. In Solovay’s method, managing shadow positions required transmitting additional $H(S)$ bits of information. We were able to reduce the number of additional bits to $O(\log H(S))$ and, thereby, improve the upper bound.
Can we reduce the upper bound even more or find a matching lower bound? In his work [5], Vereshchagin used Martin’s game to obtain an upper bound for finite sets. However, Ageev [1] proved that the quadratic bound is both a lower and upper bound in Martin’s game. Thus, to improve the upper bound further, a novel technique must be developed. It is also unknown if the lower bound in Martin’s game does extend to a lower bound of deterministic complexity.
@@ 5 Acknowledgements
We extend our gratitude to Nikolay Vereschagin, Alexander Shen, and Daniil Musatov for their invaluable assistance in both validating our findings and contributing to the writing of this article.

### 14 Dec 2023 v2

@@ Abstract
Consider a subset of positive integers $S$. In this paper, we reduce the upper bound on the length of a minimum program that enumerates $S$ in terms of the probability of $S$ being enumerated by a random program.
So far, the best-known upper bound was given by Solovay. Solovay proved that the minimum length of a program enumerating $S$ is bounded by $3$ times minus binary logarithm of the probability that a random program enumerates $S$. Later, Vereshchagin showed that the constant can be improved from $3$ to $2$ for finite sets. By improving the method proposed by Solovay, we demonstrate that any bound for finite sets implies the same bound for infinite sets, modulo logarithmic factors. Thus, the constant can be replaced by $2$ for every set $S$ due to the result of Vereshchagin.
@@
In Section 1, we introduce definitions of deterministic and randomized Kolmogorov complexity of a set. Then, we present the results known prior to our work. After that, we formulate the main result of the paper in Theorem 1.3. In Section 2, in order to prove Theorem 1.3, we describe the notions of cats and ants. Both cats and ants move along the edges of the graph $G$ described in Subsection 2.1. The position of cats at time $t$ is equal to the set enumerated by the deterministic machine halted after $t$ steps. Similarly, the position of ants at time $t$ is equal to the set enumerated by the randomized machine halted after $t$ steps. Then, we introduce shadow positions for the ants, which will follow the ants with some delay. We describe how the shadow positions will mimic the real ones. The shadow positions will correspond to an another randomized machine $M^{\prime}$, which is used in the proof of the Theorem 1.3. After that, we are finally able to state Lemma 2.1. Lastly, we prove that Theorem 1.3 follows from Lemma 2.1. In Section 3, we prove Lemma 2.1. Section 4 contains a discussion of the results.
@@ 1 Introduction
To formulate the main results of this work, we will introduce the concepts of complexities for deterministic and probabilistic enumerating machines.
@@ 1.1 Deterministic Enumerating Machine
Consider a deterministic machine $D$. $D$ takes a binary string as an input and enumerates a subset of natural numbers. $D$ is not required to halt and may enumerate an infinite set. Let us define the complexity of a set $S\subset\mathbb{N}$, denoted as $I_{D}(S)$, as the minimum length of an input for $D$ on which it enumerates $S$. If there is no such input, then $I_{D}(S)=\infty$. By a standard theorem, there exists a machine $D_{0}$ such that for any other machine $D,$ there is a constant $c$ such that:
@@
The machine $D_{0}$ is referred to as universal. Set $D$ equal to any such universal machine $D_{0}$ and define $I(S)$ as $I_{D}(S)$. Thus, $I(S)$ is well-defined up to an additive constant.
@@ 1.2 Probabilistic Enumerating Machine
Similar to the deterministic case, there exists a machine $M_{0}$ in such a way that for any other $M,$ there is a constant $c$ such that:
@@
Such a machine $M_{0}$ is referred to as optimal. Set $M$ equal to any such optimal machine $M_{0}$ and define $H(S)$ as $H_{M}(S).$ Therefore, $H(S)$ is well-defined up to an additive constant.
@@ 1.3 Connection Between Deterministic and Randomized Complexity
However, obtaining the converse inequality is more challenging. Solovay proved a linear upper bound.
Theorem 1.1.
(Solovay, [4]). There exists a constant $c$ such that for any $S\subset\mathbb{N}$, it holds that $I(S)\leq 3\cdot H(S)+2\log H(S)+c.$
Later, Vereshchagin improved the constant from three to two for finite sets:
Theorem 1.2.
(Vereshchagin, [5]). There exists a constant $c$ such that for any finite $S\subset\mathbb{N}$, it holds that $I(S)\leq 2\cdot H(S)+2\log H(S)+c.$
@@
Theorem 1.3.
Let $\alpha\geq 1$ be such that for any finite $S\subset\mathbb{N}$,
@@
Applying this Theorem 1.3 to Theorem 1.2, we improve the constant in Solovay’s bound from three to two (ignoring logarithmic factors). Thus, we obtain the following theorem.
Theorem 1.4.
There exists a constant $c$ such that for any $S\subset\mathbb{N}$, it holds that $I(S)\leq 2\cdot H(S)+O(\log H(S))+c.$
To prove Theorem 1.3, we will introduce convenient terms. We will represent the behavior of deterministic and randomized machines as the movement of cats and ants on a directed graph.
@@ 2.6 Main Lemma
Lemma 2.1.
Let $S$ be a set with a probability of enumeration greater than $\varepsilon$ (the threshold from the shift conditions). Let $S^{\prime}\subset S$ be a finite subset of it. Then at some point, a shift will occur for the set $X$ located between $S^{\prime}$ and $S$ (i.e., $S^{\prime}\subset X\subset S$).
Before proving Lemma 2.1, let’s show that it implies a desired bound for the complexity of enumerating infinite sets. Consider an increasing sequence of finite sets $S^{\prime}$ converging to $S$ and apply Lemma 2.1 to each of them, obtaining, for each, its set $X$ (and one of the first $l(\varepsilon)$ cats servicing this set). Since there is a finite number of cats, one of them must have participated in servicing an infinite number of sets $S^{\prime}$, never leaving $S$ and being in the supersets of infinitely many sets $S^{\prime}$. Therefore, it will eventually enumerate $S$. Thus, we have:
@@
Now we will show how to use this to prove Theorem 1.3.
Proof of Theorem 1.3. Let $k$ be an integer. We will show that if we set $\varepsilon=2^{-k}$ and $l(\varepsilon)=2^{\alpha k+O(\log k)}$, then during a shift, there will always be a cat among the first $l(\varepsilon)$ that comes to $X.$ Using $I(S)\leq\log_{2}l(\varepsilon),$ we will prove Theorem 1.3.
@@
$\square$
Now, it only remains to prove Lemma 2.1.
3 Proof of Lemma 2.1
@@ 3.1 Number of Shifts
At every moment $t$ of the process, the Cantor space is divided into cones corresponding to ants. For each ant, we can count the total number of shifts in which it (itself or its ancestors) participated. We can define the set $W_{k}^{t}$ as the set in the Cantor space corresponding to ants with $k$ or more shifts by the time $t$. This is a basic (open and closed) set; the larger $k$ is, the smaller $W_{k}^{t}$ (for a given $t$). On the other hand, $W_{k}^{t}$ grows with increasing $t$. We can take the union over all $t$ and obtain an open set $W_{k}$ (it is even effectively open since the process is algorithmic — we assume that $\varepsilon$ is rational). We obtain a decreasing sequence of open sets:
@@ 3.2 The Case of Non-empty Intersection
In addition to $W_{\infty}$, consider the set $U$ in the Cantor space, consisting of those sequences for which the probabilistic machine $M$ enumerates the set $S$. Do the sets $U$ and $W_{\infty}$ intersect? We will show that if they do intersect, then the statement of Lemma 2.1 is true, and then we will contradict the assumption that the intersection is empty.
@@ 4 Discussion
One of the most natural questions in theoretical computer science is how much more efficient randomized computations are compared to deterministic ones. As Theorem 1.3 demonstrates, randomized computations are not superior to deterministic ones when considering infinite sets instead of finite ones. Moreover, the gap between complexities can be reduced to a factor of 2 (modulo logarithmic factors). To prove Theorem 1.3, we utilized a trick first used by Solovay: instead of cats (the deterministic machine) trying to catch the actual positions of ants (the randomized machine), the cats catch the shadow positions of the ants, which are slightly modified real positions. In Solovay’s method, managing shadow positions required transmitting additional $H(S)$ bits of information. We were able to reduce the number of additional bits to $O(\log H(S))$ and, thereby, improve the upper bound.
Can we reduce the upper bound even more or find a matching lower bound? In [5], Vereshchagin used Martin’s game to obtain an upper bound for finite sets. However, Ageev [1] proved that the quadratic bound is both a lower and upper bound in Martin’s game. Thus, to improve the upper bound further, a novel technique must be developed. It is also unknown if the lower bound in Martin’s game does extend to a lower bound of deterministic complexity.
@@ 5 Acknowledgements
We extend our gratitude to Nikolay Vereshchagin, Alexander Shen, and Daniil Musatov for their invaluable assistance in both validating our findings and contributing to the writing of this article.